OpenStudy (anonymous):

Find the limit as x->-1 of (x^2+1)/(x+1). i've forgotten how to figure out problems when denominator = 0

OpenStudy (akshay_budhkar):

l' hospital rule, differentiate the numerator n denominator

OpenStudy (akshay_budhkar):

do that till u get a determinate form

OpenStudy (anonymous):

never learned a hospital rule

OpenStudy (anonymous):

can someone do steps pls?

OpenStudy (akshay_budhkar):

ok you differentiate the numerator and the denominator and then substitute the value of x

myininaya (myininaya):

no you can't use l'hospital here

OpenStudy (akshay_budhkar):

why :O

myininaya (myininaya):

\[(-1)^2+1=1+1=2 \neq 0\]

OpenStudy (akshay_budhkar):

lol sorry :P mathematical error :P

OpenStudy (anonymous):

can someone please just do a quick list of steps

OpenStudy (anonymous):

so i can figure out others like it

myininaya (myininaya):

the function is undefined at x=-1 this is not a hole so the limit does not exist

OpenStudy (anonymous):

yes it does, the answer is -2

myininaya (myininaya):

no the limit does not exist

myininaya (myininaya):

you typed your wrong if the answer is -2

OpenStudy (anonymous):

i did

OpenStudy (anonymous):

my fault, fml

OpenStudy (akshay_budhkar):

lol

OpenStudy (anonymous):

x^2-1****

myininaya (myininaya):

\[\lim_{x \rightarrow -1}\frac{x^2-1}{x+1}=\lim_{x \rightarrow -1}\frac{(x-1)(x+1)}{x+1}\]

OpenStudy (akshay_budhkar):

lol myininaya u r too quick :P

OpenStudy (anonymous):

damn

myininaya (myininaya):

now you also use l'hospital if you wanted to since we do have 0/0 here

myininaya (myininaya):

but i mean you don't have to you can just use some algebra

myininaya (myininaya):

like i did above (x+1)/(x+1)=1 so you are left with (x-1) just plug in -1

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

Limit is +Infinity from the right and -Infinity from the left. Plot attached.

OpenStudy (akshay_budhkar):

yeah limit at that point doesn't exist as left hand limit is not equal to right hand limit. But he made an error with the question its x^2-1 not x^+1